# Physics-Regularized Multi-Modal Image Assimilation for Brain Tumor Localization

### v1

@@ 3.2 Healthy Anatomy
A healthy brain is roughly symmetric between hemispheres, as justified by anatomical and functional observations. Imaging techniques such as MRI reveal symmetrical tissue structures, and electroencephalography (EEG) recordings show symmetrical patterns of electrical activity in the brain [26, 27]. This assumption applies to key tissue types, including white matter, gray matter, and cerebrospinal fluid. Enforcing the symmetry condition at progressively coarser resolutions aids in convergence. We construct a loss function, denoted as $L_{\text{Healthy Anatomy}}$, to quantify the asymmetry of the learned $\mathcal{F}^{G}(\mathbf{m}^{0})$. The implementation details are provided in Appendix C.
@@
Figure 3: Inference overview. (a) Patient’s MRI 3D scans. (b) Estimated tissues through non-rigid registration of the average brain, showing white matter and visible tumor segmentations. (c) Learned tumor cell distribution, regularized by the physics residual and aligned with patient data. (d) Learned initial condition of the tissues representing healthy anatomy. (e) Average brain template for reference, rigidly registered to the MRI scan. Notable anatomical differences include the lack of matter passage between the hemispheres next to the tumor, which could affect tumor cell inference results.
4 Visible Data
@@
Visible data refers to information directly observed from the patient’s MRI and FET-PET 3D scans, including visible tumor parts and surrounding brain tissues. As shown in Figure 1, this data includes estimated brain tissue probability maps, visible tumor segmentation, and a metabolic map derived from FET-PET scans, focusing on the enhancing core and edema. MRI scans provide segmentation maps [28], revealing the visible tumor. Using non-rigid atlas image registration with tumor masking [29], we estimate tissues both outside and hidden by the tumor, including white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). These observations offer partial information about both the anatomy and the tumor cell distribution.
@@
The imaging function (see Figure 2) bridges simulated tumor cell densities with visible tumor segmentation and brain tissues, as well as the metabolic map. It depends on parameters for thresholding the tumor cell distribution to obtain segmentation, denoted as $\theta_{\text{Imaging}}$, which must be learned.
@@
We construct a loss function that minimizes the difference between model predictions and patient observables. Specifically, $L_{\text{Core}}$ matches the visible tumor core, $L_{\text{Edema}}$ matches the visible edema, and $L_{\text{Metabolic Map}}$ matches the metabolic map profile. Tissue intensities carried by particles are set to reproduce the patient’s estimated tissues at $t=1$. While these intensities are predetermined, particle positions are weakly constrained, so matching the tissues (WM, GM, CSF) to the observed structures outside the tumor is enforced through $L_{\text{Tissues}}$. The constructed penalty terms align our spatiotemporal distributions of tumor cells and tissues with patient data characteristics. Implementation details of the data losses and the imaging function are provided in Appendix B.
@@ 5 Combined Loss Function
The loss function $\mathcal{L}_{\text{Total}}$ is constructed to balance the contributions of physics-based regularization, assumptions about the initial conditions, and the visible empirical data. This facilitates a comprehensive approach to infer the full spatial tumor cell distribution at the time of the patient’s data acquisition (see Figure 3 for an overview of the inference process and Figure 2 for the learning process visualization):
@@
$\,\underbrace{\alpha_{6}L_{\text{Core}}+\alpha_{7}L_{\text{Edema}}+\alpha_{8}L_{\text{Metabolic Map}}+\alpha_{9}L_{\text{Tissues}}}_{\text{Visible Data}}$ (10)
All $\alpha_{*}$ parameters were selected through synthetic experiments with both single focal and local multi-focal tumors, as explained in Appendix D.
@@ 6 Validation and Ablation Study of Radiotherapy Planning
It is important to note that there is no objective ground truth for the tumor cell distribution of pre-operative tumors. The performances of methods are evaluated through the efficacy at a downstream task of predicting tumor recurrence present at a follow-up MRI scan within a volume defined by inferred tumor cell concentration  [^2] The Data-Driven Neural Networks (Unconstrained) is trained end-to-end to predict tumor recurrences from which we extract the radiotherapy volume around it, similar to the distance map (Appendix A). isosurfaces. The total volume of the region encompassed by the isosurfaces adheres to the current clinical practice, which uses a 1.5 cm distance isosurface around the pre-operative MRI-segmented tumor core, referred to as the Standard Plan, excluding any areas outside the brain tissues. Figure 4 visualizes the Standard Plan and a plan based on our learned tumor cell distribution. Each model’s target volume is simply a redistribution of the Standard Plan volume.
Our evaluation metric is Recurrence Coverage [%], which measures the percentage of the tumor segmentation, as detected on follow-up MRIs, that is encompassed within the radiotherapy target volume defined by each model. Additionally, follow-up MRI scans showing tumor recurrence are registered to the pre-operative anatomy using rigid registration, ensuring that predictions and actual recurrences are aligned accurately.
Figure 4: Radiotherapy planning. a) Pre-operative segmentations with white matter concentration as background. b) Distance map from the tumor core segmentation with a 1.5 cm contour and within diffusive tissue, constituting the Standard Plan. c) Our learned tumor cell distribution with the isosurface contour where the enclosed volume equals the total volume of the Standard Plan.
@@ 6.2 Dataset
We use a publicly available dataset of 58 patients with preoperative MRI scans and preoperative FET-PET imaging, along with follow-up MRI scans at the first visible tumor recurrence [16] (MIT licence).
@@ 6.3 Models and Features
  • Data-Driven Neural Networks (Physics-Constained: Similar to the above unconstrained approach but limited to finding parameters for physics simulations, incorporating hard physics constraints.
@@ 6.4 Results
Table 1, using the "Recurrence Coverage [%] (Enhancing Core)" column, and Figure 5b show results where recurrence is defined by enhancing core segmentation, per RANO guidelines [31]. Our method leads with 89.9% average recurrence coverage, compared to 89.02% for static grid discretization and 87.3% for the Standard Plan. Numerical physics simulations are more robust at 86.2% compared to 84.31% for the physics-constrained data-driven method. The unconstrained data-driven method remains the lowest at 66.8%. Results are close because the enhancing core often occurs near pre-operative MRI segmentations, leading to many methods achieving 100% coverage (seen in ’Equal’ in Figure 5b). However, Figure 5b shows an increase in the preference from 19% to 28%.
@@
Figure 5: Direct patient-by-patient comparisons to the Standard Plan of radiotherapy plans with equal total volumes. "Greater," "Equal," and "Less" refer to the direct numerical comparison of Recurrence Coverage. a) Recurrence is defined as the union of edema, enhancing core, and necrotic core on the follow-up MRI segmentation. b) Recurrence is defined as the enhancing core on the follow-up MRI segmentation.
@@ Appendix A Baselines Implementation
Numerical Physics Simulations
@@
: We employ the numerical forward simulation scheme from [35] which was implemented on GPU [7]. The simulations follows the equation 1. We initialize the brain tissues with a rigidly registered average brain. For $\theta_{\text{Dynamics}}$, $x_{0}\in\theta_{\text{Initial}}$, and $\theta_{\text{Imaging}}$, we use the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) method [5, 36].
Data-Driven Neural Networks (Unconstrained)
@@
: We use nnU-Net [37] which takes brain tissue distribution, FET-PET map, and pre-operative tumor segmentations to directly predict core tumor recurrences. After predicting the recurrences, we construct a distance distribution from the predicted recurrences. This distribution is then thresholded, similar to other methods, to create a binary radiotherapy treatment map of the same volume as the Standard Plan. The network is trained using an 80%/20% train-test split.
Data-Driven Neural Networks (Physics Constrained)
@@
: We employ a recently published method [7] that uses CNNs to map pre-operative tumor segmentation to $\theta_{\text{Dynamics}}$ and runs a numerical simulation following equation 1. This method uses an average brain as the initial condition. After inference, the tumor cell distribution is non-rigidly registered to the patient’s brain anatomy.
Static Grid Discretization
@@
: This method, as implemented in [16], uses estimated brain tissues (4) as the initial condition for the anatomy and later keeps the tissues static, effectively setting $\gamma\in\theta_{\text{Dynamics}}$ in Equation 3 to 0.
Standard Plan
@@
: This approach uses a 1.5 cm uniform safety margin around the resection cavity and/or remaining tumor on MRI imaging, aligning with both North American and European guidelines [38, 39].
Appendix B Imaging Model
@@
$L_{\text{Visible Data}}=\alpha_{6}L_{\text{Core}}(c,\theta_{\text{up}})+\alpha_{7}L_{\text{Edema}}(c,\theta_{\text{down}},\theta_{\text{up}})+\alpha_{8}L_{\text{Metabolic Map}}(c)+\alpha_{9}L_{\text{Tissues}}(\boldsymbol{m})$ (11)
@@
$\sigma(x)=\frac{1}{1+e^{-\beta x}}$ (12)
@@
$L_{\text{Core}}(c,\theta_{\text{up}})=\sum_{i=0}^{N_{x}}\sigma(\theta_{\text{up}}-c^{N_{t}}_{i}-\alpha)$ (13)
@@
$L_{\text{Edema}}(c,\theta_{\text{down}},\theta_{\text{up}})=\sum_{i=0}^{N_{x}}\sigma(\theta_{\text{down}}-u^{N_{t}}_{i}-\alpha)+(1-\sigma(\theta_{\text{up}}-c^{N_{t}}_{i}+\alpha))$ (14)
@@
In this context, $\Omega_{2}$ represents the collection of voxels that map to the time point at which the imaging is conducted; for single-image analysis, this corresponds to the final time slice.
@@
$L_{\text{PET}}(c)=1-\text{corr}(c^{N_{t}},p^{\text{PET}})$ (15)
Here, $\text{corr}(u^{n},p^{\text{PET}})$ represents the Pearson correlation coefficient between the predicted tumor cell densities $c^{N_{t}}$ and the restricted FET-PET signal $p^{\text{PET}}$, across all voxels within within the edema and enhancing core regions, providing a direct measure of how well the model predictions align with the observed metabolic profiles.
Appendix C Assumption of Brain Symmetry at Initial Time Without the Tumor

### v3

@@ 3.2 Healthy Anatomy
The deformations are clearly visible in the MRI images and are an integral part of understanding growth dynamics. To obtain a plausible estimate of the unperturbed brain, we need to incorporate an anatomical prior.
@@
A healthy brain is roughly symmetric between hemispheres, as justified by anatomical and functional observations. Imaging techniques such as MRI reveal symmetrical tissue structures, and electroencephalography (EEG) recordings show symmetrical patterns of electrical activity in the brain [32, 33]. This assumption applies to key tissue types, including white matter, gray matter, and cerebrospinal fluid. We construct a loss function, denoted as $L_{\text{Healthy Anatomy}}$, to quantify the asymmetry of the learned $\mathcal{F}^{G}(\mathbf{m}^{0})$. The implementation details are provided in Appendix E. We do not use the unperturbed anatomy explicitly for inference of the tumor cells, however, we need additional constraints to bound $\gamma$ from Eq. 3 that couples tumor cells with tissue dynamics. While symmetry of the healthy brain is a coarse assumption, we soften it by quantifying the asymmetry in lower resolutions.
@@
Figure 3: Inference overview: (a) Patient’s MRI 3D scans. (b) Estimated tissues through non-rigid registration of the average brain, showing white matter and visible tumor segmentations. (c) Learned tumor cell distribution, regularized by the physics residual and aligned with patient data. (d) Learned initial condition of the tissues representing healthy anatomy. (e) Average brain template for reference, rigidly registered to the MRI scan. Notable anatomical differences include the lack of matter passage between the hemispheres next to the tumor, which could affect tumor cell inference results.
4 Imaging Data Reconstruction
@@
Imaging data refers to information directly observed from the patient’s MRI and FET-PET 3D scans [34], including visible tumor parts and surrounding brain tissues. As shown in Figure 1, this data includes estimated brain tissue probability maps, visible tumor segmentation, and a metabolic map derived from FET-PET scans, focusing on the enhancing core and edema. MRI scans provide segmentation maps [35], revealing the visible tumor. Using non-rigid atlas image registration with tumor masking [34], we estimate tissues both outside and hidden by the tumor, including white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). These observations offer partial information about both the anatomy and the tumor cell distribution.
@@
The imaging function (see Figure 2) [22, 36] bridges simulated tumor cell densities with visible tumor segmentation and brain tissues, as well as the metabolic map. It depends on parameters for thresholding the tumor cell distribution to obtain segmentation, denoted as $\theta_{\text{Imaging}}$, which must be learned.
@@
We construct a loss function that minimizes the difference between model predictions and patient observables. Specifically, $L_{\text{Core}}$ matches the visible tumor core, $L_{\text{Edema}}$ matches the visible edema, and $L_{\text{Metabolic Map}}$ matches the metabolic map profile. Tissue intensities carried by particles are set to reproduce the patient’s estimated tissues at $t=1$. While these intensities are predetermined, particle positions are weakly constrained, so matching the tissues (WM, GM, CSF) to the observed structures outside the tumor is enforced through $L_{\text{Tissues}}$. The constructed penalty terms align our spatiotemporal distributions of tumor cells and tissues with patient data characteristics. Implementation details of the data losses and the imaging function are provided in Appendix D.
@@ 5 Combined Loss Function
The loss function $\mathcal{L}_{\text{Total}}$ is constructed to balance the contributions of physics-based regularization, assumptions about the initial conditions, and the imaging empirical data. This facilitates a comprehensive approach to infer the full spatial tumor cell distribution at the time of the patient’s data acquisition (see Figure 3 for an overview of the inference process and Figure 2 for the learning process visualization):
@@
$\,\underbrace{\alpha_{5}L_{\text{Core}}+\alpha_{6}L_{\text{Edema}}+\alpha_{7}L_{\text{Metabolic Map}}+\alpha_{8}L_{\text{Tissues}}}_{\text{Imaging Data Reconstruction}}.$ (10)
All $\alpha_{*}$ parameters were determined based on experiments with synthetic data involving both single focal and local multi-focal tumors. Extensive ablation studies are provided in Appendix C, and the details of the synthetic experiment setup can be found in Appendix F.
@@ 6 Validation and Ablation Study of Radiotherapy Planning
There is no objective ground truth for pre-operative tumor cell distribution, so methods are evaluated through downstream tasks like defining radiation target volumes. An "accurate" method targets areas that later show progression in follow-up MRI, acknowledging that radiation delays but does not eliminate tumor regrowth. An "inaccurate" method spares these areas or targets regions that remain tumor-free. The current "Standard Plan" defines the target volume using a 1.5 cm isosurface around the MRI-segmented tumor core, while our approach generates isosurfaces from learned tumor cell distributions. Figure 4 visualizes the Standard Plan and a plan based on our learned tumor cell distribution. Each model’s target volume is simply a redistribution of the target volume of the Standard Plan.
We assess accuracy by measuring the overlap between the target volumes and regions of later tumor progression, calculating the Recurrence Coverage [%]. This metric represents the percentage of the tumor progression region that falls within the radiotherapy target volume of a given model, after rigid alignment of both scans.
Additional experiments studying properties of the tumor growth and imaging model on synthetic data, i.e., without relying on indirect validation via radiotherapy planning, are given in Appendix B.
Figure 4: Radiotherapy planning: a) Pre-operative segmentations with white matter concentration as background. b) Distance map from the tumor core segmentation with a 1.5 cm contour and within diffusive tissue, constituting the Standard Plan. c) Our learned tumor cell distribution with the isosurface contour where the enclosed volume equals the total volume of the Standard Plan.
@@ 6.2 Dataset
We use a publicly available dataset of 58 patients with preoperative MRI scans and preoperative FET-PET imaging, along with follow-up MRI scans at the time of the first visible tumor recurrence [22].
We would like to emphasize that this is the largest publicly available dataset of this kind, i.e., with multi-modal imaging data and follow-up MRI. Its size is common for medical studies and sufficient to motivate clinical trials (see, for example, [38, 39, 40] on clinical investigations into the use of PET for radiotherapy planning). The problem we address is a critical challenge in cancer treatment research that calls for ML solutions. However, data acquisition is costly and time-consuming, and clinical datasets, such as ours with only a few dozen complete observations, are insufficient to support purely data-driven ML approaches (as visible from results reported in Table 1). At the same time, it is this inherent limitation that is encouraging us to incorporate physics-based assumptions and physics-informed regularizations into an ML model to effectively address the driving clinical problem.
The follow-up images with visible tumor recurrences are not used by our method during the optimization or hyper-parameter search at any point; therefore, the dataset can be considered 100% unseen. The decision not to use any recurrence data for hyper-parameter tuning is to avoid diluting the already small dataset.
@@ 6.3 Models and Features
  • Data-Driven Neural Networks (Physics-Constrained): Similar to the above unconstrained approach but limited to finding parameters for physics simulations, incorporating hard physics constraints.
@@ 6.4 Results
Table 1, using the "Recurrence Coverage [%] (Enhancing Core)" column, and Figure 5b show results where recurrence is defined by enhancing core segmentation, per RANO guidelines [41]. Our method leads with 89.9% average recurrence coverage, compared to 89.0% for static grid discretization and 87.3% for the Standard Plan. Numerical physics simulations are more robust at 86.2% compared to 84.3% for the physics-constrained data-driven method. The unconstrained data-driven method remains the lowest at 66.8%. Results are close because the enhancing core often occurs near pre-operative MRI segmentations, leading to many methods achieving 100% coverage (seen in ’Equal’ in Figure 5b). However, Figure 5b shows an increase in the preference from 19% to 28%.
@@
Figure 5: Direct patient-by-patient comparisons to the Standard Plan of radiotherapy plans with equal total volumes: "Greater," "Equal," and "Less" refer to the direct numerical comparison of Recurrence Coverage. a) Recurrence is defined as the union of edema, enhancing core, and necrotic core on the follow-up MRI segmentation. b) Recurrence is defined as the enhancing core on the follow-up MRI segmentation.
As shown in Figure 5, the gap between the “Greater” and “Less” categories represents the number of patients who benefit from our method compared to the standard plan. We observe a consistent improvement with our method, demonstrating superior performance beyond just higher mean recurrence coverage, which could be skewed by outliers. The results are statistically significant, as shown in Table 3 in Appendix B.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
We extend our gratitude to John Lowengrub and Ray Zirui Zhang from the University of California, Irvine, for their invaluable insights and discussions. This research was supported by the Helmut Horten Foundation and the DCoMEX project.
@@ Appendix A Baselines Implementation
Numerical Physics Simulations:
@@
We employ the numerical forward simulation scheme from [45] which was implemented on GPU [13]. The simulations follows the equation 1. We initialize the brain tissues with a rigidly registered average brain. For $\theta_{\text{Dynamics}}$, $x_{0}\in\theta_{\text{Initial}}$, and $\theta_{\text{Imaging}}$, we use the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) method [10, 46].
Data-Driven Neural Networks (Unconstrained):
@@
We use nnU-Net [47] which takes brain tissue distribution, FET-PET map, and pre-operative tumor segmentations to directly predict core tumor recurrences. After predicting the recurrences, we construct a distance distribution from the predicted recurrences. This distribution is then thresholded, similar to other methods, to create a binary radiotherapy treatment map of the same volume as the Standard Plan. The network is trained using an 80%/20% train-test split.
Data-Driven Neural Networks (Physics-Constrained):
@@
We employ a recently published method [13] that uses CNNs to map pre-operative tumor segmentation to $\theta_{\text{Dynamics}}$ and runs a numerical simulation following equation 1. This method uses an average brain as the initial anatomy. After inference, the tumor cell distribution is non-rigidly registered to the patient’s brain anatomy.
Static Grid Discretization:
@@
This method, as implemented in [22], uses estimated brain tissues (4) as the initial condition for the anatomy and later keeps the tissues static, effectively setting $\gamma\in\theta_{\text{Dynamics}}$ in Equation 3 to 0.
Standard Plan:
@@
This approach uses a 1.5 cm uniform safety margin around the resection cavity and/or remaining tumor on MRI imaging, aligning with both North American and European guidelines [48, 49].
Appendix B Synthetic Results
To show how the method performs on tasks where there is ground truth, our method and GliODIL [16] (previous SOTA, Static Grid) are additionally evaluated on synthetic cases generated by numerical physics solvers under various conditions. See Table 2 for the results and Table 3 for statistical tests.
RMSE (1 Focal Origin) [%] RMSE (3 Focal Origins) [%]
Input Type Ours Static Grid [16] Ours Static Grid [16]
Tumor Core $\textbf{10.2}\pm 0.4$ $12.2\pm 0.6$ $13.3\pm 0.5$ $\textbf{13.0}\pm 0.7$
Tumor Core, Edema, and Metabolic Map $\textbf{2.5}\pm 0.3$ $5.5\pm 0.8$ $\textbf{4.0}\pm 0.5$ $6.9\pm 0.5$
Tumor Core, Edema, and Metabolic Map,
$\theta_{\text{Initial}}$, $\theta_{\text{Dynamics}}$
$\textbf{0.3}\pm 0.1$ $3.3\pm 0.2$ $\textbf{0.4}\pm 0.1$ $4.2\pm 0.3$
Table 2: Synthetic results comparing RMSE for 30 patients with different input types for one and three tumor origins, using our method and Static Grid Discretization (GliODIL [16]). The test dataset was not used for hyper-parameter tuning. Parameter ranges in Appendix D and $\gamma\in[0,1.5]$ from Eq. 3 in the manuscript. Ours wins because GliODIL cannot deform brain tissues through its static grid limitations.
Comparison Recurrence (Any) Recurrence (Enhancing Core)
p-value p-value
Ours vs Standard Plan (1.5cm) 0.014 0.00082
Ours vs Static Grid Discretization (GliODIL[16]) 0.043 0.039
Table 3: Statistical significance tests comparing recurrence coverage (any and enhancing core) among different models using the Wilcoxon signed-rank test. The data is not normally distributed which we tested using Shapiro-Wilk test resulting in p-values below 1E-12.
Appendix C Extended Ablation Study
Additional information regarding the loss coefficients:
$\alpha_{2}$ is essential for grid stability. Without it, the grid would be unstable, causing folding and rendering the results meaningless. If set to zero, $\alpha_{2}$ can be replaced by a Laplacian-based regularization, as visualized in Figure 6b and Figure 6c, though this clearly results in less biologically plausible deformations.
$\alpha_{4}$ regularizes the initial anatomy to bound $\gamma$ from Eq 3. It should be kept low due to the natural asymmetry in brains. More details are provided in Appendix E.
$\alpha_{5}$, $\alpha_{6}$, $\alpha_{7}$, and $\alpha_{8}$ are crucial for grounding PDE solutions in reality (most important, all other losses can be seen simply as regularisation of these components). Imaging data reconstruction loss cannot be calibrated using synthetic studies as all the other alphas. Instead of using recurrence data to calibrate them, we set them such that they provide a balanced contribution to the total loss
$\alpha_{3}$ should be kept low since the initial tumor shape is unknown. However, decreasing it too much causes a local minimum with no tumor.
[…]
Removing elements of the imaging data reconstruction loss introduces ill-posedness and degradation of synthetic results (see Table 4). Removing tumor growth physics regularizations effectively reduces our solution to performance close to the standard plan. Removing particle dynamics (static grid) reduces our model to [16] Static Grid Discretization (“GliODIL”).
Figure 6: Deformation through particle movement: a) Total loss of our model for reference. b) Particle position with tissue elasticity. c) Particle position without tissue elasticity. These particle positions are learned for the pre-operative time. For grid stability in c), additional regularization is required. For $\alpha_{2}>0$ we observe one large localized deformation caused by the tumor, which is more biologically plausible than multiple smaller deformations visible for $\alpha_{2}=0$.
Model Variation Recurrence (Any) Recurrence (Enhancing Core)
Coverage [%] IoU* Coverage [%] IoU*
Ours, $\alpha_{5}=0,\alpha_{6}=0,\alpha_{7}=0,\alpha_{8}=0$ $60.5\pm 4.9$ $0.15\pm 0.02$ $83.5\pm 3.7$ $0.05\pm 0.01$
Ours, $\alpha_{6}=0,\alpha_{7}=0$ $63.5\pm 4.0$ $0.15\pm 0.03$ $86.5\pm 3.6$ $0.06\pm 0.02$
Ours, $\alpha_{3}=0$ $70.0\pm 3.8$ $0.16\pm 0.02$ $87.4\pm 3.4$ $0.06\pm 0.02$
Standard Plan (1.5cm) $70.0\pm 3.8$ $0.16\pm 0.02$ $87.3\pm 3.6$ $0.04\pm 0.01$
Ours, $\alpha_{1}=0$ $70.9\pm 3.8$ $0.16\pm 0.02$ $87.6\pm 3.4$ $0.05\pm 0.02$
Ours, $\alpha_{2}=0$ $71.9\pm 3.8$ $0.15\pm 0.02$ $88.1\pm 3.4$ $0.05\pm 0.02$
Ours, $\alpha_{4}=0$ $72.8\pm 3.9$ $0.16\pm 0.02$ $89.0\pm 3.4$ $0.06\pm 0.02$
Static Grid Discretization (GliODIL, [16]) $72.9\pm 3.5$ $0.16\pm 0.02$ $89.0\pm 3.3$ $0.06\pm 0.02$
Ours $\textbf{74.7}\pm 3.1$ $\textbf{0.19}\pm 0.02$ $\textbf{89.9}\pm 2.7$ $\textbf{0.08}\pm 0.02$
Table 4: Ablation study showing the impact of different model variations on recurrence coverage and IoU. Large deviations account for various sizes of tumors and recurrences, yet the results remain consistent and statistically significant (Table 3). *IoU heavily penalizes the fact that the radiotherapy volumes are equal between baselines and much larger than typical recurrences. Equal volume is to make comparisons with the Standard Plan without volume/recall trade-offs.
Appendix D Imaging Model
@@
$L_{\text{Imaging Data Reconstruction}}=\alpha_{5}L_{\text{Core}}(c,\theta_{\text{up}})+\alpha_{6}L_{\text{Edema}}(c,\theta_{\text{down}},\theta_{\text{up}})+\alpha_{7}L_{\text{Metabolic Map}}(c)+\alpha_{8}L_{\text{Tissues}}(\boldsymbol{m}).$ (11)
@@
$\sigma(x)=\frac{1}{1+e^{-\beta x}}.$ (12)
@@
$L_{\text{Core}}(c,\theta_{\text{up}})=\sum_{i=0}^{N_{x}}\sigma(\theta_{\text{up}}-c^{N_{t}}_{i}-\alpha),$ (13)
@@
$L_{\text{Edema}}(c,\theta_{\text{down}},\theta_{\text{up}})=\sum_{i=0}^{N_{x}}\sigma(\theta_{\text{down}}-c^{N_{t}}_{i}-\alpha)+(1-\sigma(\theta_{\text{up}}-c^{N_{t}}_{i}+\alpha)),$ (14)
@@
$L_{\text{PET}}(c)=1-\text{corr}(c^{N_{t}},p^{\text{PET}}).$ (15)
Here, $\text{corr}(c^{N_{t}},p^{\text{PET}})$ represents the Pearson correlation coefficient between the predicted tumor cell densities $c^{N_{t}}$ and the restricted FET-PET signal $p^{\text{PET}}$, across all voxels within within the edema and enhancing core regions, providing a direct measure of how well the model predictions align with the observed metabolic profiles.
Appendix E Assumption of Brain Symmetry at Initial Time Without the Tumor